Solve these differential equations:

1.) dy/dt = P[(1/y)-1] +by-a

2.) dy/dt = b*y* e^(-ct)

'P', 'b', and 'a' are constants.

Thanks.
Answer #1

(1)

gdy dt e [ 3 - 1] + by a = P(1-4) g by² - ay-Pytp a. + by - y by? - Catply + 1 dt. y dy by? (apply + P y dy b yz atpy , dt. -

(2)

bé et dt. 2) d = bye -et b, c are constant. on dy , ỹ on, & bath sides. on taking integration 6 feet s dy dt. -ct be on lnya